# Causal progress with imperfect placebo treatments and outcomes

### v2

@@ Abstract
In the quest to make defensible causal claims from observational data, it is sometimes possible to leverage information from “placebo treatments” and “placebo outcomes”. Existing approaches employing such information focus largely on point identification and assume (i) “perfect placebos”, meaning placebo treatments have precisely zero effect on the outcome and the real treatment has precisely zero effect on a placebo outcome; and (ii) “equiconfounding”, meaning that the treatment-outcome relationship where one is a placebo suffers the same amount of confounding as does the real treatment-outcome relationship, on some scale. We instead consider an omitted variable bias framework, in which users can postulate ranges of values for the degree of unequal confounding and the degree of placebo imperfection. Once postulated, these assumptions identify or bound the linear estimates of treatment effects. Our approach also does not require using both a placebo treatment and placebo outcome, as some others do. While applicable in many settings, one ubiquitous use-case for this approach is to employ pre-treatment outcomes as (perfect) placebo outcomes, as in difference-in-difference. The parallel trends assumption in this setting is identical to the equiconfounding assumption, on a particular scale, which our framework allows the user to relax. Finally, we demonstrate the use of our framework with two applications and a simulation, employing an R package that implements these approaches.
@@ 1 Introduction
Unobserved confounding plagues researchers seeking to draw causal inferences from observational data. When we cannot rule out unobserved confounding, we may look to alternative strategies to fully or partially identify the range of defensible causal effects given the assumptions we can support. One set of approaches rests on observing variables that (i) may be affected by similar factors as those affecting the treatment, but which cannot affect the outcome (placebo treatments), and/or (ii) variables sharing similar influences with the outcome, but that cannot be affected by treatment (placebo outcomes). [^2] In related literatures such as the proximal causal inference literature discussed more below, the terms “negative control outcome” and “negative control exposure” have been used to reference “placebo outcome” and “placebo treatment”, respectively. The broad intuition is then that, if these variables are indeed similar to their real counterparts in terms of the degree of confounding they suffer, they allow us to estimate and remove that much confounding from the real treatment-outcome relationship.
@@
To validate and demonstrate these tools, we provide a simulated example and two real-world applications. First, we create a simple simulation focused on the effect of regular physical activity on incidence of heart disease with a placebo treatment. Next, we use data from the well-known National Supported Work Demonstration to show that under reasonable assumptions an investigator given the observational data would have been able to produce a range of estimates that contains the experimental benchmark value. In an additional application, we show that despite apparent evidence to the contrary, the Zika virus likely decreased birth rates in Brazil in 2016. We conduct these analyses using the PlaceboLM package, which we make available in the R computing language.
@@ 2.1 Causal structures of interest
Placebo treatments and outcomes, perfect or imperfect, are defined by the role they play in some causal model. We employ directed acyclic graphs (DAGs) (Pearl, 1995) to label these roles. Figure 1 shows simple structures with “perfect” placebo treatments ($P$) and placebo outcomes ($N$). Placebo treatments and outcomes can look very similar in their causal role; indeed Figure 1(a) and (b) are mirror images, replacing the role of $P$ with $N$. We explore such ambiguities in the context of a broader set of cases provided in Section  2.6. To call a placebo “imperfect” is to allow for an additional causal effect, and so always implies an additional edge in the causal structure. For example, Figure 1(c) adds the $D\to N$ edge to Figure 1(a), while Figure 1(d) adds the $P\to Y$ edge to Figure 1(b).
@@
2.2 An omitted variable bias approach to placebos: Placebo outcomes
@@
The key maneuver for making this useful is to then rewrite the bias on the coefficient of interest due to the omission of $\mathbf{Z}$ as a function of the bias in the treatment-placebo relationship. A first way to do so is simply to define $m\in\mathbb{R}$ such that:
@@
The value of $m$ is not known in reality, but the supposition will be that we can postulate a range of values for it and see the consequence. For any postulated value of $m$, we can solve for $\hat{\beta}_{Y\sim D|Z,X}$ at that value of $m$ by algebraic rearragnement,
@@ 2.5 Placebo treatments
Similar machinery can leverage imperfect placebo treatments rather than imperfect placebo outcomes. Suppose that we want to run the long regression for $Y$ on $D$ (Equation 1) but, again, $\mathbf{Z}$ is an unobserved vector of variables that we would have liked to include in the regression but cannot. Suppose also that we have a placebo treatment, $P$, for which we believe the confounding of $P$ and $Y$ from $\mathbf{Z}$ can be compared to the confounding of $D$ and $Y$ from $\mathbf{Z}$. We might consider running two separate regressions as we did for placebo outcomes. However, that approach is not easily adapted to imperfect placebo treatments, where there may be a direct relationship between $P$ and $Y$ [^8] Suppose we have a causal model like Figure 1(d). $\hat{\beta}_{Y\sim Z|P,X}$ (or $\hat{R}_{Y\sim Z|P,X}$) will capture the causal relationship $D\to Y$, since the path $Z\to D\to Y$ will remain open conditional on $P$ and $X$ alone. Expressing the bias in the regression coefficient on $D$ from the short regression using a two regression approach like we used for placebo outcomes would, therefore, require us to reason about the causal effect of interest itself, which is not useful. Instead, suppose that $P$ is not a descendant of $D$, $D$ is not a descendant of $P$, $P$ is a good or neutral control for the effect of $D$ on $Y$, and $P$ is a good or neutral control for $D$ on $Y$, conditional on $\mathbf{Z}$ and $\mathbf{X}$ [^9] We require that $D$ and $P$ are good or neutral controls for each other in that they do not increase the bias in our estimate of the coefficient we are targeting ($\hat{\beta}_{Y\sim D|X,Z}$). This will simplify interpretation. If a user is concerned that there could be a path opened or closed by conditioning on $D$ or $P$, we suggest drawing the DAG to check. See Section 2.6 for further discussion. Further, suppose that we want to run the “long” regression of $Y$ on $D,P,\mathbf{X}$, and $\mathbf{Z}$ in Equation 17. In particular, we are interested in the coefficient $\beta_{Y\sim D|P,Z,X}$. However, since $\mathbf{Z}$ is unobserved, we must run the “short” regression of $Y$ on $D,P$ and $\mathbf{X}$ in Equation 18, rather than the desired long regression.
@@ 2.6 Varieties of (single) placebos
2.7 Difference in difference as a special case
@@ 3.1 Simulation evidence with a placebo treatment
To first demonstrate our procedure and verify that it behaves as expected we construct a simple simulation for a plausible example involving a placebo treatment. Suppose we are interested in the effect that physical activity has on cardiovascular health for adults in the United States. We have a representative sample of adults with data on whether they exercise regularly or not ($D$), whether they take supplemental vitamins ($P$), and whether they experience heart disease over a fixed follow up period ($Y$). We may worry that socioeconomic status ($Z$), a variable we do not observe in our data, confounds the observed exercise–heart disease ($D\to Y$) relationship. For example, socioeconomic status could influence whether or not individuals are able to exercise regularly while influencing risk of heart disease through numerous other channels (diet, access to healthcare, quality of living conditions, stress, etc). We also suspect that such factors could influence whether or not individuals demonstrate other health seeking behaviors, such as supplemental vitamins. Further, suppose that we expect these supplements to have little or no effect on heart disease. Thus, supplemental vitamins ($P$) can be used as a (possibly imperfect) placebo treatment in our effort to determine the causal effect of regular exercise on heart disease.
@@
3.2 National Supported Work Demonstration
@@
3.3 Effects of Zika virus on birth rates
In a second application, we investigate the effect Zika virus had on birth rates in Brazil. Zika virus can cause birth complications, which may lead either to couples delaying planned pregnancies, or to increased risk of miscarriage. We use data previously analyzed by Taddeo et al. (2022) and Tchetgen et al. (2023), also employing placebo-based techniques, but requiring different assumptions. In Brazil in 2015, the state of Pernambuco saw high rates of Zika, while zero cases were reported in the state of Rio Grande do Sul. As in previous works, we regard the former state as the treated group and the later as the control group. We use municipality level overall birth rates per 1000 people in 2016 as our primary outcome of interest, while the same measure in 2014 is our placebo outcome. We make the reasonable assumption that the treatment has no direct causal effect on the pre-treatment (placebo) outcome. Using data provided by Amorim (2022), we perform the analysis using the PlaceboLM package; see Appendix F for code.
@@ 4 Discussion
While this approach applies in settings that do not resemble DID, the connection to DID is notable and relates it to a growing literature and range of ways in which investigators might consider relaxing the parallel trends assumption. Prior work has also noted the connection between placebo outcomes and DID (Sofer et al., 2016). The parallel trends assumption is frequently difficult to support: if units may select into treatment on the basis of unobserved factors that impact the outcomes, it is difficult to then argue that no unobserved factor related to treatment assignment influences the average trend in the (non-treatment) outcome. Thus, additional tools that allow reasoned relaxations of the parallel trends assumptions are highly valuable. Numerous other approaches to relaxing the parallel trends assumption, or to improve on the current practice of testing for parallel trends in the pre-treatment period have been proposed (see e.g. Manski and Pepper, 2018; Bilinski and Hatfield, 2018; Freyaldenhoven et al., 2019; Keele et al., 2019; Ryan et al., 2019; Ye et al., 2020; Gibson and Zimmerman, 2021; Rambachan and Roth, 2023). The principal difference between any of these approaches and ours (when applied to DID settings) lies in what they ask the user to reason about. For example, Rambachan and Roth (2023) suggest that users consider violations of parallel trends to have magnitudes that are multiples of how different the treated and control group were on trends entirely prior to treatment. Gibson and Zimmerman (2021) applies the logic of omitted variable bias as in Cinelli and Hazlett (2020) to DID, by using a first-difference regression when data are panel structured. This asks users to reason about how much of (i) the overtime change in the (non-treatment) and (ii) the overtime change in treatment activation are explained by unobserved confounding. Our approach requires investigators to reason about how much bias is suffered in the post-treatment outcome difference, as a multiple of the bias suffered in a placebo (pre-treatment) outcome difference (i.e. $m$, or $k$). Our approach also does not require that the placebo outcome is the same variable or on the same scale, as required for pre-treatment outcomes in DID. Which of these approaches offers users the greatest leverage is a question of the specific domain and the information available. Our view is that whichever approach facilitates understanding and determining plausible bounds in a specific context is to be preferred. For example, investigators may be better equipped to argue that the confounding experienced by the pre-treatment outcome is between 0.5 and 2 times the confounding experienced by the post-treatment outcome in a context where they are not able to bound the differential trends in simple additive terms. By contrast, if the trends in the outcome absent treatment can be constrained by domain knowledge or physical limitations, then arguments about the additive difference in non-treatment trends may be more useful. Further, we argue that any approach that shows results as a function of an unknown “identification parameter” (like $m$ or $k$) can be valuable and transparency-improving, simply because it reminds the reader that conventional DID estimates stand on a heroic assumption. That is, while users may argue they cannot defend any particular value of $m$ or $k$ in our setting, simply ignoring this fact by asserting that $m=1$ (as in conventional DID) is not credibility-enhancing.
@@
Table 1: Taxonomy of Single Placebo Graphs
Graph ($\mathbf{X}$ omitted for clarity) Placebo Type Short Regression(s) Parameters Target Coefficient Expression
[…]
[b] $D$$Y$$P$$\mathbf{Z}$ Placebo Outcome (see above row); $\beta_{P\sim D|X,Z}\neq 0$
[c] $D$$Y$$P$$\mathbf{Z}$ Placebo Treatment (see above row); $\beta_{Y\sim P|D,Z,X}\neq 0$
[…]
[d] $D$$Y$$P$$\mathbf{Z}$ See Zhang and Ding (2022) for partial identification of indirect and direct effects
Could use Cinelli and Hazlett (2020) for partial identification of total effects; $R^{2}_{Y\sim Z|D,X}$ includes $Z\to P\to Y$
Placebo Outcome (see above row); $\beta_{P\sim D|X,Z}$ is part of total effect; $\text{bias}_{\text{(YD.X)}}$ includes $Z\to P\to Y$
[…]
Table 2: Taxonomy of Single Placebo Graphs - Continued
Graph ($\mathbf{X}$ omitted for clarity) Placebo Type Short Regression(s) Parameters Target Coefficient Expression
[…]
[f] $D$$Y$$P$$\mathbf{Z}$ Observed Confounder 2 (see above row)
[…]
[h] $D$$Y$$P$$\mathbf{Z}$ Post-Outcome (see above row)
Appendix

### v3

@@ Abstract
In the quest to make defensible causal claims from observational data, researchers may leverage information from “placebo treatments” and “placebo outcomes”. Existing approaches assume (i) perfect placebos (meaning placebo treatments have no effect on outcomes and the real treatment has no effect on placebo outcomes) and (ii) equiconfounding (equal confounding across placebo and real treatment-outcome relationships). Many also require the availability of a valid placebo treatment and placebo outcome. We propose a new framework based on omitted variable bias, allowing users to postulate ranges of values corresponding to unequal confounding and imperfect placebos. These assumptions identify or bound linear treatment effect estimates, with either a placebo outcome or treatment. While applicable in many settings, one case for this method is the use of pre-treatment outcomes as placebo outcomes, as in difference-in-difference. In this setting, the parallel trends assumption is identical to equiconfounding on a specific scale, to which our framework provides a reasoned relaxation. We demonstrate this approach in applications that examine the effect of a job training program, and another that estimates the effect of the Zika virus on birth rates in Brazil. These analyses employ the R package placeboLM package.
Keywords: Causal inference, omitted variable bias, proximal inference, placebos, partial identification
@@ 1 Introduction
Across disciplines, unobserved confounding plagues researchers seeking to draw causal inferences from observational data. When we cannot rule out unobserved confounding, we may look to alternative strategies to fully or partially identify the range of defensible causal effects given the assumptions we can support. One set of approaches rests on observing variables that (i) may be affected by similar factors as those affecting the treatment, but which cannot affect the outcome (placebo treatments), and/or (ii) variables sharing similar influences with the outcome, but that cannot be affected by treatment (placebo outcomes). [^2] In related literatures such as the proximal causal inference literature discussed more below, the terms “negative control outcome” and “negative control exposure” have been used to reference “placebo outcome” and “placebo treatment”, respectively. The broad intuition is then that, if these variables are indeed similar to their real counterparts in terms of the degree of confounding they suffer, they allow us to estimate and remove that much confounding from the real treatment-outcome relationship.
@@
After a simulation example, we apply this approach to two real-world investigations. First, we use data from the well-known National Supported Work Demonstration to show that under reasonable assumptions an investigator given the observational data would have been able to produce a range of estimates that contains the experimental benchmark value. Second, we show that despite apparent evidence to the contrary, the Zika virus likely decreased birth rates in Brazil in 2016. We conduct these analyses using the PlaceboLM package, which we make available in the R computing language.
@@ 2.1 Causal structures of interest
Placebo treatments and outcomes, perfect or imperfect, are defined by the role they play in some causal model. We employ directed acyclic graphs (DAGs) (Pearl, 1995) to label these roles. Figure 1 shows simple structures with “perfect” placebo treatments ($P$) and placebo outcomes ($N$). Placebo treatments and outcomes can look very similar in their causal role; indeed Figure 1(a) and (b) are mirror images, replacing the role of $P$ with $N$. We explore such ambiguities in the context of a broader set of cases in the Online Supplement. To call a placebo “imperfect” is to allow for an additional causal effect, and so always implies an additional edge in the causal structure. For example, Figure 1(c) adds the $D\to N$ edge to Figure 1(a), while Figure 1(d) adds the $P\to Y$ edge to Figure 1(b).
@@
2.2 An omitted variable bias approach to placebo outcomes
@@
The key maneuver for making this useful is to then rewrite the bias on the coefficient of interest due to the omission of $\mathbf{Z}$ as a function of the bias in the treatment-placebo relationship. A first way to do so is simply to define $m\in\mathbb{R}$ such that
@@
The value of $m$ is not known in reality, but the supposition will be that we can postulate a range of values for it and see the consequence. For any postulated value of $m$, we can solve for $\hat{\beta}_{Y\sim D|Z,X}$ at that value of $m$ by algebraic rearrangement,
@@ 2.5 Placebo treatments
Similar machinery can leverage imperfect placebo treatments rather than imperfect placebo outcomes. Suppose that we want to run the long regression for $Y$ on $D$ (Equation 1) but, again, $\mathbf{Z}$ is an unobserved vector of variables that we would have liked to include in the regression but cannot. Suppose also that we have a placebo treatment, $P$, for which we believe the confounding of $P$ and $Y$ from $\mathbf{Z}$ can be compared to the confounding of $D$ and $Y$ from $\mathbf{Z}$. We might consider running two separate regressions as we did for placebo outcomes. However, that approach is not easily adapted to imperfect placebo treatments, where there may be a direct relationship between $P$ and $Y$ [^8] Suppose we have a causal model like Figure 1(d). $\hat{\beta}_{Y\sim Z|P,X}$ (or $\hat{R}_{Y\sim Z|P,X}$) will capture the causal relationship $D\to Y$, since the path $Z\to D\to Y$ will remain open conditional on $P$ and $X$ alone. Expressing the bias in the regression coefficient on $D$ from the short regression using a two regression approach like we used for placebo outcomes would, therefore, require us to reason about the causal effect of interest itself, which is not useful. Instead, suppose that $P$ is not a descendant of $D$, $D$ is not a descendant of $P$, $P$ is a good or neutral control for the effect of $D$ on $Y$, and $P$ is a good or neutral control for $D$ on $Y$, conditional on $\mathbf{Z}$ and $\mathbf{X}$ [^9] We require that $D$ and $P$ are good or neutral controls for each other in that they do not increase the bias in our estimate of the coefficient we are targeting ($\hat{\beta}_{Y\sim D|X,Z}$). This will simplify interpretation. If a user is concerned that there could be a path opened or closed by conditioning on $D$ or $P$, we suggest drawing the DAG to check. See the Supplement for further discussion. Further, suppose that we want to run the “long” regression of $Y$ on $D,P,\mathbf{X}$, and $\mathbf{Z}$ in Equation 17. In particular, we are interested in the coefficient $\beta_{Y\sim D|P,Z,X}$. However, since $\mathbf{Z}$ is unobserved, we must run the “short” regression of $Y$ on $D,P$ and $\mathbf{X}$ in Equation 18, rather than the desired long regression,
@@ 2.6 Varieties of (single) placebos
2.7 Relaxed difference in difference as a special case
@@ 3.1 Simulation evidence with a placebo treatment
Before applying the approach, we demonstrate it and verify that it behaves as expected in a simple simulation representing a plausible placebo treatment example. Suppose we are interested in the effect that physical activity has on cardiovascular health for adults in the United States. We have a representative sample of adults with data on whether they exercise regularly or not ($D$), whether they take supplemental vitamins ($P$), and whether they experience heart disease over a fixed follow up period ($Y$). We may worry that socioeconomic status ($Z$), a variable we do not observe in our data, confounds the observed exercise–heart disease ($D\to Y$) relationship. For example, socioeconomic status could influence whether or not individuals are able to exercise regularly while influencing risk of heart disease through numerous other channels (diet, access to healthcare, quality of living conditions, stress, etc). We also suspect that such factors could influence whether or not individuals demonstrate other health seeking behaviors, such as supplemental vitamins. Further, suppose that we expect these supplements to have little or no effect on heart disease. Thus, supplemental vitamins ($P$) can be used as a (possibly imperfect) placebo treatment in our effort to determine the causal effect of regular exercise on heart disease.
@@
3.2 Estimating the effect of a job training program on earnings
@@
3.3 Estimating the effects of Zika virus on birth rates in Brazil
In a second application, we investigate the effect Zika virus had on birth rates in Brazil. Zika virus can cause birth complications, which may lead either to couples delaying planned pregnancies, or to increased risk of miscarriage. We use data previously analyzed by Taddeo et al. (2022) and Tchetgen et al. (2023), also employing placebo-based techniques, but requiring different assumptions. In Brazil in 2015, the state of Pernambuco saw high rates of Zika, while zero cases were reported in the state of Rio Grande do Sul. As in previous works, we regard the former state as the treated group and the later as the control group. We use municipality level overall birth rates per 1000 people in 2016 as our primary outcome of interest, while the same measure in 2014 is our placebo outcome.
We make the reasonable assumption that the treatment has no direct causal effect on the pre-treatment (placebo) outcome. Using data provided by Amorim (2022), we perform the analysis using the PlaceboLM package; see Appendix F for code.
@@ 4 Discussion
While this approach applies in settings that do not resemble DID, the connection to DID is notable and relates it to a growing literature and range of ways in which investigators might consider relaxing the parallel trends assumption. Prior work has also noted the connection between placebo outcomes and DID, such as (Sofer et al., 2016), which also proposes a method to employ placebo outcomes on different scales from the true outcome. Scale aside, the parallel trends assumption is frequently difficult to support, however: if units may select into treatment on the basis of unobserved factors that impact the outcomes, it is difficult to then argue that no unobserved factor related to treatment assignment influences the average trend in the (non-treatment) outcome. Thus, additional tools that allow reasoned relaxations of the parallel trends assumptions are highly valuable. Numerous other approaches to relaxing the parallel trends assumption, or to improve on the current practice of testing for parallel trends in the pre-treatment period have been proposed (see e.g. Manski and Pepper, 2018; Bilinski and Hatfield, 2018; Freyaldenhoven et al., 2019; Keele et al., 2019; Ryan et al., 2019; Ye et al., 2020; Gibson and Zimmerman, 2021; Rambachan and Roth, 2023). The principal difference between any of these approaches and ours (when applied to DID settings) lies in what they ask the user to reason about. For example, Rambachan and Roth (2023) suggest that users consider violations of parallel trends to have magnitudes that are multiples of how different the treated and control group were on trends entirely prior to treatment. Gibson and Zimmerman (2021) applies the logic of omitted variable bias as in Cinelli and Hazlett (2020) to DID, by using a first-difference regression when data are panel structured. This asks users to reason about how much of (i) the overtime change in the (non-treatment) and (ii) the overtime change in treatment activation are explained by unobserved confounding. Our approach requires investigators to reason about how much bias is suffered in the post-treatment outcome difference, as a multiple of the bias suffered in a placebo (pre-treatment) outcome difference (i.e. $m$, or $k$). Our approach also does not require that the placebo outcome is the same variable or on the same scale, as required for pre-treatment outcomes in DID. Which of these approaches offers users the greatest leverage is a question of the specific domain and the information available. Our view is that whichever approach facilitates understanding and determining plausible bounds in a specific context is to be preferred. For example, investigators may be better equipped to argue that the confounding experienced by the pre-treatment outcome is between 0.5 and 2 times the confounding experienced by the post-treatment outcome in a context where they are not able to bound the differential trends in simple additive terms. By contrast, if the trends in the outcome absent treatment can be constrained by domain knowledge or physical limitations, then arguments about the additive difference in non-treatment trends may be more useful. Further, we argue that any approach that shows results as a function of an unknown “identification parameter” (like $m$ or $k$) can be valuable and transparency-improving, simply because it reminds the reader that conventional DID estimates stand on a heroic assumption. That is, while users may argue they cannot defend any particular value of $m$ or $k$ in our setting, simply ignoring this fact by asserting that $m=1$ (as in conventional DID) is not credibility-enhancing.